OpenStudy (anonymous):

Find the Range of \[\tfrac{x^2 + 3x + 5}{x^2 + x-6}\]

OpenStudy (anonymous):

(x^2 +3x + 5) / (x^2 + x - 6)

OpenStudy (experimentx):

try factorizing the denominator

OpenStudy (experimentx):

x^2 + x - 6 = (x+3)(x-2) check for range in this interval (-inf, -3),(-3,2),(2, inf)

OpenStudy (anonymous):

then...)

OpenStudy (experimentx):

what is the min value and max value of the expression interval (-inf, -3)?

OpenStudy (anonymous):

y = (x^2 +3x + 5) / (x^2 + x - 6) y(x^2 + x - 6)=(x^2 +3x + 5)

OpenStudy (anonymous):

x^2 y + xy -6y = x^2 + 3x + 5 x^2 (y-1) + x( y-3) - (6y + 5) = 0 (y-3)^2 + 4 (y-1) (6y+5) >=0

OpenStudy (anonymous):

is this correct.) @experimentX

OpenStudy (anonymous):

y^2 - 6y + 9 + 4 ( 6y^2 + 5y - 6y - 5) > =0 25y^2 - 10 y -11 > =0

OpenStudy (anonymous):

and i am stuck here......)

OpenStudy (experimentx):

x^2 (y-1) + x( y-3) - (6y + 5) = 0 x is always real ... so (y-3)^2 + 4(y-1)(6y+5) >= 0

OpenStudy (anonymous):

Yup....i have taken it...)

OpenStudy (anonymous):

25y^2 - 10 y -11 > =0 ............ i cant factorize this....)

OpenStudy (experimentx):

complete the square ... not sure this will work. http://www.wolframalpha.com/input/?i=25y^2+-+10+y+-11+%3E%3D+0

OpenStudy (anonymous):

ok....) thxxx.....

OpenStudy (experimentx):

seems you did your work correctly.